OpenStudy (anonymous):

If a line with a slope of -2 crosses the y-axis at (0, 3), what is the equation of the line?

OpenStudy (anonymous):

y − 2x = 3 -2y + 3x = 1 y + 3x = 2 y + 2x = 3

OpenStudy (anonymous):

@iGreen

OpenStudy (texaschic101):

y = mx + b slope(m) = -2 (0,3)...x = 0 and y = 3 now we sub 3 = -2(0) + b 3 = b so your equation is : y = -2x + 3 but we need it in standard form Ax + By = C y = -2x + 3 -- add 2x to both sides 2x + y = 3 (answer D)

OpenStudy (igreen):

y = mx + b Where m = slope(in this case -2), and where b = the y-intercept(in this case 3)

OpenStudy (anonymous):

Thanks!!!!! But now dont know who to give the medal..